# Exact mobility edges in quasiperiodic network models with slowly varying potentials

### v2

@@ Abstract
Quasiperiodic models are important physical platforms to explore Anderson transitions in low dimensional systems, yet the exact mobility edges (MEs) are generally hard to be determined analytically. To date, the MEs in only a few models can be determined exactly. In this manuscript, we propose a new class of network models characterized by quasiperiodic slowly varying potentials and the absence of hidden self-duality, and exactly determine their MEs. We take the mosaic models with slowly varying potentials as examples to illustrate this result and derive its MEs from the effective Hamiltonian. In this method, we can integrate out the periodic sites to obtain an effective Hamiltonian with energy-dependent potentials $g(E)V$ and effective eigenenergy $f(E)$, which directly yields the MEs at $f(E)=\pm(2t^{\kappa}\pm g(E)V)$, where $\kappa\in\mathbb{Z}^{+}$. With this idea in hand, we then generalize our method to more quasiperiodic network models, including those with much more complicated geometries and non-Hermitian features. Finally, we propose the realization of these models using optical waveguides and show that the Anderson transition can be observed even in small physical systems (with lattice sites about $L=50-100$). Our results provide some key insights into the understanding and realization of exact MEs in experiments.
@@ I Introduction
Disorder is always an important subject of interest in condensed matter physics. The study of localization phenomena in disordered low-dimensional systems has been one of the central problems in condensed matter physics [1, 2, 3, 4]. In one-dimensional (1D) systems, the scaling theory of localization indicates that any weak random potential leads all states to be localized [5, 6, 7, 8, 9, 10]. However, in models with correlated disorders, such as those involving quasiperiodic potentials [11, 12, 13, 14], Anderson transitions occur at some finite critical strengths. Thus, 1D models with quasiperiodic potentials can exhibit much rich physics, including localized, extended, and critical phases, along with the associated mobility edges (MEs) [15, 16, 17, 18]. A pedagogical example of such a quasiperiodic system is the Aubry-André-Harper (AAH) model [11, 19, 20, 21, 22, 23, 24], in which the self-duality between real space and momentum space leads to the exact phase boundaries separating localized and extended states at some critical potential strength $V_{\rm c}=\pm 2t$ [11]. Consequently, all states are extended when $|V|<|V_{c}|$, while they become localized for $|V|>|V_{c}|$, with all states at the critical points exhibit multifractality [25, 26, 27]. In Ref. [28], we show that this potential in the quasiperiodic network models can directly yield energy-dependent MEs, and in Ref. [chen2025reentrantba], we investigate the Anderson transition in the quasiperiodic model using optical waveguides. Furthermore, the Anderson transitions in 1D models have been intensively explored in experiments with ultracold atoms [29, 30, 31], acoustic systems [32, 33] and photonic systems [34, 35, 36].
Figure 1: (a) The 1D mosaic model with a slowly varying potential for $\kappa=2$. (b) The network model when $N=3$, which can be interpreted as a multipath mosaic model, where $N$ denotes the number of paths. More network models can be found in Ref. [28]. (c) The effective quasiperiodic chain after integrating out the periodic sites. Here, $V_{i}$ represents the quasiperiodic slowly varying potential, for example $V\cos(\pi\alpha i^{\nu}+\phi)$, and $gV_{j}$ is the corresponding effective potential.
@@
can be exactly solved using semianalytical technique [57, 58, 59, 60, 61, 62]. In particular, Das Sarma et al. demonstrated that the 1D quasiperiodic models with this potential can support MEs at $E_{c}=\pm(2t-V)$ for $10<\nu<$ 1 and $V<2$, while for $\nu>1$, all states away from the exact band center are localized [59].
@@
Recently, we demonstrated a method for determining the MEs of various quasiperiodic network models with self-dual potentials [28], which is constituted by periodic network sites and quasiperiodic sites. We show that although the original models lack self-duality during Fourier transformation to the momentum space, the effective models obtained by integrating out the periodic sites exhibit a hidden self-duality. This method naturally yields energy-dependent MEs. As compared to the above models, the quasiperiodic network models offer some significant flexibility, enabling us to design the MEs in more physical models. The major idea of this work is to extend the discussion of MEs from network models with self-dual potentials in Ref. [28] to those with slowly varying potentials that lack self-duality. In this manuscript, we theoretically obtain the exact MEs in a new class of 1D quasiperiodic network models with slowly varying potentials. We take the potential in Eq. (1) as an example to illustrate this idea. In our method, by integrating out the periodic sites, we obtain an effective tight-binding model, from which the exact MEs can be determined in terms of the energy-dependent potential $g(E)V$ and effective eigenenergy $f(E)$. Following the semianalytical method proposed by previous work [59], we can directly obtain the MEs at $f(E)=\pm(2t\pm g(E)V)$. In this method, $g(E)=0$ naturally yields the resonant states, which separate the extended states from the localized states. Finally, we propose the physical realization of these network models using optical waveguides and demonstrate that their associates Anderson transition in principle can be observed using some small physical systems, with lattice sites about $L=50-100$. In this way, we expect the physics models proposed in this work can be realized in the near future in experiments.
@@
Figure 2: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for mosaic models with slowly varying potentials when (a) $\kappa=2$ and (b) $\kappa=3$. The solid lines denote the MEs given by Eqs. (6), (7), and (8), respectively. (c)-(e) Spatial distributions of three representative eigenstates with $E\approx-0.55t$ (extended), $E\approx-0.45t$ (localized), and $E=0$ (resonant) as marked in Fig. 2(a). The other parameters are $\alpha=(\sqrt{5}-1)/2$, $\nu=0.6$, $\phi=0$ and $L=987$.
@@
The rest of the manuscript is organized as follows. Section II introduces the Hamiltonian of mosaic models with slowly varying potentials and presents a method to solve their MEs. Next, in Sec. III, we generalize our results to network models with slowly varying potentials and obtain their MEs. In Sec. IV, we propose to realize these models using optical waveguides and demonstrate that the transition from extended phase to localized phase may be observed even with small physical systems ($L=50-100$), which is currently realizable in experiments [63]. Finally, we summarize and discuss these results in Sec. V.
@@ II.1 Hamiltonian and MEs
where $c_{i}^{\dagger}$ ($c_{i}$) represents the creation (annihilation) operator at site $i$, and $t$ denotes the nearest-neighbor hopping coefficient. For simplicity, we set $t=1$ as the energy unit. The on-site energy $V_{i}$ follows a mosaic pattern and is defined as
$V_{i}$ =$=\left\{\begin{array}[]{lr}V{\rm cos}(\pi\alpha i^{\nu}+\phi),\hskip 8.5359pt\kappa|i\\ 0,\hskip 71.98543pt\kappa\nmid i\end{array}\right..$ (5)
@@
We can gain some insight into the physics of this model by considering some special limits, which have already been studied in literature. When $\nu=1$ and $\kappa=1$, the model reduces to the well-known AAH model, which exhibits a phase transition between extended and localized states. Specifically, all states are extended for $|V|<2t$ and localized for $|V|>2t$, with $V=\pm 2t$ being the self-dual points where all states are critical [11]. For $\nu=1$ and $\kappa\neq 1$, the model becomes the quasiperiodic mosaic AAH model, where the hidden self-duality ensures the presence of MEs, such as $E_{c}=\pm 2t^{2}/V$ for the minimal case $\kappa=2$, and $E_{c}=\pm\sqrt{t^{2}\pm 2t^{3}/V}$ for $\kappa=3$ [41]. When $10<\nu<$ 1 and $\kappa=1$, the model corresponds to a quasiperiodic model with slowly varying potential, where extended states are located at the central region of the band ($|E|<2t-V$) and localized states appear at the band edge ($|E|>2t-V$). For $\nu>1$ and $\kappa=1$, all states away from the exact band center are localized [59].
In this manuscript we focus on the condition that $10<\nu<$ 1 and $\kappa\neq 1$. One of the central results of this study is that the MEs in this model can be exactly determined from the effective Hamiltonian, obtained by integrating out the periodic sites. For the minimal case of $\kappa=2$, there are four exact energy-dependent MEs given by
@@
with $V_{\text{max}}=2t$ for $E=0$ and $V_{\text{max}}=t$ for $E=\pm\sqrt{3}t$. In this case, two resonant states exist at $E=\pm t$. In the following, we first present the numerical results in Sec. II.2, followed by detailed derivations in Sec. II.3.
@@ II.2 Numerical results
We can characterize the localization properties of the wave functions by calculating their inverse participation ratio (IPR). For a normalized wave function |$\ket{\psi^{n}}=\sum_{i}u_{i}^{n}c_{i}^{\dagger}\ket{0}$, the IPR is defined as
${\rm IPR^{n}}=\sum_{i=1}^{L}|u_{i}^{n}|^{4},$ (9)
where $n$ represents the band index. For an extended eigenstate, with $u_{i}^{n}\sim 1/\sqrt{L}$, the IPR scales as ${\rm IPR}^{n}\propto L^{-1}$, which approaches zero in the thermodynamic limit. Conversely, for a localized state, the IPR maintains a finite value as the system size increases, thus ${\rm IPR}^{n}\propto\mathcal{O}(1)$. For a critical state, the IPR scales as ${\rm IPR}^{n}\propto L^{-D_{2}}$, where the fractal dimension $D_{2}\in(0,1)$. To distinguish the extended, localized, and critical states, we can define the fractal dimension as
$D_{2}^{n}=-{\ln{\rm IPR^{n}}\over\ln L}.$ (10)
This number resembles the Hausdorff dimension defined in fractal geometries. This number can be used to characterize these different phases when and only when $L$ is large enough.
@@
Figure 2(a) plots the fractal dimension of the corresponding eigenstates as a function of the energy $E$ and potential strength $V$ for $\kappa=2$. The black lines in the figure represent the MEs given by Eq. (6). Consistent with the analytical results, the fractal dimension $D_{2}$ is approximately one for energies satisfying $V<|E|<(\sqrt{V^{2}+16t^{2}}-V)/2$ and zero for other energies. For any $\kappa$, the model exhibits $2\kappa$ MEs, similar to the mosaic AAH model [41]. The validity of the MEs is further confirmed by the spatial distributions of three typical eigenstates, marked by pink squares in Fig. 2(a). Figures 2(c)-2(e) display the spatial distribution of these eigenstates, corresponding to $E\approx-0.55t$ (extended), $-0.45t$ (localized), and 00 (resonant) when $V=0.5t$, respectively. These results are consistent with Eq. (6), which predicts extended states for $0.5t<|E|<1.76556t$. Additionally, we plot the fractal dimension of the corresponding eigenstates for $\kappa=3$ in Fig. 2(b), which agrees well with the MEs given by Eqs. (7) and (8).
@@
where $\Lambda_{i}$ represents the eigenvalues of the matrix $\mathbf{\Lambda}$. In this case, the smallest Lyapunov exponent corresponds to the condition with longest localization length by definition $\gamma=1/\xi$. Therefore, for the localized we expect $\gamma$ is finite; while for the extended phase, we expect that $\gamma L>1$ (or $>>1$), where $L$ is the chain length. We solve this matrix using the QR decomposition method. In Fig. 3, we plot the Lyapunov exponent $\gamma$ versus $L$ of the quasiperiodic slowly varying model for various $V$ when $E=0$, showing that this value is convergent, which is necessary for a well-defined exponent. These results may also indicate ergodicity, which is necessary during the calculation of the above expression.
@@
To further investigate the localization properties of the quasiperiodic mosaic model with slowly varying potentials in the thermodynamic limit, we plot the Lyapunov exponent $\gamma$ as a function of energy $E$ for various $V$ in Fig. 4. As the energy varies, $\gamma$ changes dramatically from a finite value (indicating localized states) to 0 (indicating extended states), or vice versa, at certain energies. This abrupt jumping in $\gamma$ suggests the presence of MEs in the energy spectrum. When $\kappa=2$, we observe that the Lyapunov exponent at the exact band center is vanishingly small ($\ll 1/L$) for different values of $V$ [see Fig. 4(a)], indicating a resonant state. The MEs are clearly observed around $E_{c}\approx\pm 1.8t$ and ±$\pm 0.5t$ for $V=0.5t$, and around $E_{c}\approx\pm 1.4t$ for $V=\sqrt{2}t$, which closely align with the exact expression of MEs given by Eq. (6). When $\kappa=3$, we can also find the Lyapunov exponent $\gamma$ for states satisfying $E=\pm t$ is always zero, indicating two resonant states [see Fig. 4(b)]. Additionally, the regime with $\gamma=0$ is consistent with Eqs. (7) and (8).
@@ II.3 Analytical solutions
${\bf u}=\begin{bmatrix}\vdots\\ u_{i-3}\\ u_{i-2}\\ u_{i-1}\\ u_{i}\\ u_{i+1}\\ u_{i+2}\\ u_{i+3}\\ \vdots\end{bmatrix}.$ (19)
From the equation for $u_{i-1}$ and $u_{i+1}$, where the potentials are zero, we can obtain
@@
Figure 6: Function $\cos(\pi\alpha j^{\nu})$ with (a) $\nu=1.6$ and (b) $\nu=0.6$. (c) Function $s=(E-c)E$ with $c>0$ (left panel) and $c<0$ (right panel), and their corresponding regime for $0<s<4t^{2}$, with $t=1$. The red dots denote the MEs.
Next we provide a more detailed explanation of the above results using the method developed by Das Sarma et al. [59]. The effective potential $\lambda_{j}$ exhibits the property of becoming locally constant in the thermodynamic limit
@@
due to $10<\nu<$ 1 [see Fig. 6(b)]. In contrast when $\nu>1$, the oscillating of the potential will become more and more rapid, yielding fully localized states in the thermodynamic limit [see Fig. 6(a)].
@@
with $-V\leq c\leq V$, yielding the following simplified equation
@@
The above results have an immediate corollary, that is, the phase boundaries are independent of the phase carried in the potential, assuming $V_{j}=V\cos(\pi\alpha j^{\nu}+\phi)$. The reason is that while this phase will influence the potential at each site, it will not influence the maximal and minimal value of the potential. Meanwhile, the parameters of $\alpha$ is also not important, nor it is not necessary to be an irrational number. We find that $\alpha$ and $\phi$ will never appear in the expression of MEs. These are also verified by our numerical simulation, which are not shown.

### v3

@@ Abstract
Quasiperiodic models are important physical platforms to explore Anderson transitions in low dimensional systems, yet the exact mobility edges (MEs) are generally hard to be determined analytically. To date, the MEs in only a few models can be determined exactly. In this manuscript, we propose a new class of network models characterized by quasiperiodic slowly varying potentials and the absence of hidden self-duality, and exactly determine their MEs. We take the mosaic models with slowly varying potentials as examples to illustrate this result and derive its MEs from the effective Hamiltonian. In this method, we can integrate out the periodic sites to obtain an effective Hamiltonian with energy-dependent potentials $g(E)V$ and effective eigenenergy $f(E)$, which directly yields the MEs at $f(E)=\pm(2t^{\kappa}\pm g(E)V)$, where $\kappa\in\mathbb{Z}^{+}$. With this idea in hand, we then generalize our method to more quasiperiodic network models, including those with much more complicated geometries and non-Hermitian features. Finally, we propose the realization of these models using optical waveguides and show that the Anderson transition can be observed even in small physical systems (with lattice sites about L=50100L=50-100). Our results provide some key insights into the understanding and realization of exact MEs in experiments.
@@ I Introduction
Disorder is always an important subject of interest in condensed matter physics. The study of localization phenomena in disordered low-dimensional systems has been one of the central problems in condensed matter physics [1, 2, 3, 4]. In one-dimensional (1D) systems, the scaling theory of localization indicates that any weak random potential leads all states to be localized [5, 6, 7, 8, 9, 10]. However, in models with correlated disorders, such as those involving quasiperiodic potentials [11, 12, 13, 14], Anderson transitions occur at some finite critical strengths. Thus, 1D models with quasiperiodic potentials can exhibit much rich physics, including localized, extended, and critical phases, along with the associated mobility edges (MEs) [15, 16, 17, 18]. A pedagogical example of such a quasiperiodic system is the Aubry-André-Harper (AAH) model [11, 19, 20, 21, 22, 23], in which the self-duality between real space and momentum space leads to the exact phase boundaries separating localized and extended states at some critical potential strength $V_{\rm c}=\pm 2t$ [11]. Consequently, all states are extended when $|V|<|V_{c}|$, while they become localized for $|V|>|V_{c}|$, with all states at the critical points exhibit multifractality [24, 25, 26]. In Ref. [27], we show that this potential in the quasiperiodic network models can directly yield energy-dependent MEs, and in Ref. [28], we investigate the Anderson transition in the quasiperiodic model using optical waveguides. Furthermore, the Anderson transitions in 1D models have been intensively explored in experiments with ultracold atoms [29, 30, 31], acoustic systems [32, 33] and photonic systems [34, 35, 36].
Figure 1: (a) The 1D mosaic model with a slowly varying potential for $\kappa=2$. (b) The network model when $N=4$, which can be interpreted as a multipath mosaic model, where NN denotes the number of paths. More network models can be found in Ref. [27]. (c) The effective quasiperiodic chain after integrating out the periodic sites. Here, $V_{i}$ represents the quasiperiodic slowly varying potential, for example $V\cos(\pi\alpha i^{\nu}+\phi)$, and $gV_{j}$ is the corresponding effective potential.
@@
can be exactly solved using semianalytical technique [57, 58, 59, 60, 61, 62]. In particular, Das Sarma et al. demonstrated that the 1D quasiperiodic models with this potential can support MEs at $E_{c}=\pm(2t-V)$ for $0<\nu<1$ and $V<2$, while for $\nu>1$, all states away from the exact band center are localized [59].
@@
Recently, we demonstrated a method for determining the MEs of various quasiperiodic network models with self-dual potentials [27], which is constituted by periodic network sites and quasiperiodic sites. We show that although the original models lack self-duality during Fourier transformation to the momentum space, the effective models obtained by integrating out the periodic sites exhibit a hidden self-duality. This method naturally yields energy-dependent MEs. As compared to the above models, the quasiperiodic network models offer some significant flexibility, enabling us to design the MEs in more physical models. The major idea of this work is to extend the discussion of MEs from network models with self-dual potentials in Ref. [27] to those with slowly varying potentials that lack self-duality. We theoretically obtain the exact MEs in a new class of 1D quasiperiodic network models with slowly varying potentials. We take the potential in Eq. (1) as an example to illustrate this idea. In our method, by integrating out the periodic sites, we obtain an effective tight-binding model, from which the exact MEs can be determined in terms of the energy-dependent potential $g(E)V$ and effective eigenenergy $f(E)$. Following the semianalytical method proposed by previous work [59], we can directly obtain the MEs at $f(E)=\pm(2t\pm g(E)V)$. In this method, $g(E)=0$ naturally yields the resonant states, which separate the extended states from the localized states. Finally, we propose the physical realization of these network models using optical waveguides and demonstrate that their associates Anderson transition in principle can be observed using some small physical systems, with lattice sites about L=50100L=50-100. In this way, we expect the physics models proposed in this work can be realized in the near future in experiments.
@@
Figure 2: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for mosaic models with slowly varying potentials when (a) $\kappa=2$ and (b) $\kappa=3$. The solid lines denote the MEs given by Eqs. (6), (7), and (8), respectively. (c)-(e) Spatial distributions of three representative eigenstates with $E\approx-0.67t$ (extended), $E\approx-0.38t$ (localized), and $E=0$ (resonant) as marked in Fig. 2(a). The other parameters are $\alpha=(\sqrt{5}-1)/2$, $\nu=0.6$, $\phi=0$ and $L=987$.
@@
The rest of the manuscript is organized as follows. Section II introduces the Hamiltonian of mosaic models with slowly varying potentials and presents a method to solve their MEs. Next, in Sec. III, we generalize our results to network models with slowly varying potentials and obtain their MEs. In Sec. IV, we propose to realize these models using optical waveguides and demonstrate that the transition from extended phase to localized phase may be observed even with small physical systems (L=50100L=50-100), which is currently realizable in experiments [28]. Finally, we summarize and discuss these results in Sec. V.
@@ II.1 Hamiltonian and MEs
where $c_{i}^{\dagger}$ ($c_{i}$) represents the creation (annihilation) operator at site $i$, and $t$ denotes the nearest-neighbor hopping coefficient. For simplicity, we set $t=1$ as the energy unit. The on-site energy $V_{i}$ follows a mosaic pattern and is defined as [63]
$V_{i}$ $=\left\{\begin{array}[]{lr}V{\rm cos}(\pi\alpha i^{\nu}+\phi),\hskip 8.5359pt{\rm mod}(i,\kappa)=0,\\ 0,\hskip 71.98543pt{\rm otherwise},\end{array}\right$. (5)
@@
We can gain some insight into the physics of this model by considering some special limits, which have already been studied in literature. When $\nu=1$ and $\kappa=1$, the model reduces to the well-known AAH model, which exhibits a phase transition between extended and localized states. Specifically, all states are extended for $|V|<2t$ and localized for $|V|>2t$, with $V=\pm 2t$ being the self-dual points where all states are critical [11]. For $\nu=1$ and $\kappa\neq 1$, the model becomes the quasiperiodic mosaic AAH model, where the hidden self-duality ensures the presence of MEs, such as $E_{c}=\pm 2t^{2}/V$ for the minimal case $\kappa=2$, and $E_{c}=\pm\sqrt{t^{2}\pm 2t^{3}/V}$ for $\kappa=3$ [41]. When $0<\nu<1$ and $\kappa=1$, the model corresponds to a quasiperiodic model with slowly varying potential, where extended states are located at the central region of the band (|E|<2tV|E|<2t-V) and localized states appear at the band edge (|E|>2tV|E|>2t-V). For $\nu>1$ and $\kappa=1$, all states away from the exact band center are localized [59].
In this manuscript we focus on the condition that $0<\nu<1$ and $\kappa\neq 1$. We show that the MEs in this model can be exactly determined from the effective Hamiltonian, obtained by integrating out the periodic sites. For the minimal case of $\kappa=2$, there are four exact energy-dependent MEs given by
@@
with $V_{\text{max}}=2t$ for $E=0$ and $V_{\text{max}}=t$ for $E=\pm\sqrt{3}t$. In this case, two resonant states exist at $E=\pm t$. These results may also be obtained using the transfer matrix method for the 1D Mosaic model in Fig. 1(a) [63]. In the following, we first present the numerical results in Sec. II.2, followed by detailed derivations in Sec. II.3.
@@ II.2 Numerical results
[…]
${\rm IPR^{n}}=\sum_{i=1}^{L}|u_{i}^{n}|^{4},\quad{\rm NPR}^{n}=\frac{1}{L}\frac{1}{\sum_{i=1}^{L}|u_{i}^{n}|^{4}},$ (9)
where $n$ represents the band index. For an extended eigenstate, with $u_{i}^{n}\sim 1/\sqrt{L}$, the IPR scales as ${\rm IPR}^{n}\propto L^{-1}$, which approaches zero in the thermodynamic limit. In this case, ${\rm NPR}^{n}\sim\mathcal{O}(1)$. Conversely, for a localized state, the IPR maintains a finite value as the system size increases, thus ${\rm IPR}^{n}\propto\mathcal{O}(1)$, thus ${\rm NPR}^{n}\sim 1/L$. For a critical state, the IPR scales as ${\rm IPR}^{n}\propto L^{-D_{2}}$, where the fractal dimension $D_{2}\in(0,1)$. To distinguish the extended, localized, and critical states, we can define the fractal dimension as
[…]
This number resembles the Hausdorff dimension defined in fractal geometries. This number has be used to characterize these different phases when and only when $L$ is large enough [52].
@@
Figure 2(a) plots the fractal dimension of the corresponding eigenstates as a function of the energy $E$ and potential strength $V$ for $\kappa=2$. The black lines in the figure represent the MEs given by Eq. (6). Consistent with the analytical results, the fractal dimension $D_{2}$ is approximately one for energies satisfying $V<|E|<(\sqrt{V^{2}+16t^{2}}-V)/2$ and zero for other energies. For any $\kappa$, the model exhibits $2\kappa$ MEs, similar to the mosaic AAH model [41]. The validity of the MEs is further confirmed by the spatial distributions of three typical eigenstates, marked by pink squares in Fig. 2(a). Figures 2(c)-2(e) display the spatial distribution of these eigenstates, corresponding to $E\approx-0.55t$ (extended), 0.45t-0.45t (localized), and 0 (resonant) when $V=0.5t$, respectively. These results are consistent with Eq. (6), which predicts extended states for $0.5t<|E|<1.76556t$. Additionally, we plot the fractal dimension of the corresponding eigenstates for $\kappa=3$ in Fig. 2(b), which agrees well with the MEs given by Eqs. (7) and (8).
@@
where $\Lambda_{i}$ represents the eigenvalues of the matrix $\mathbf{\Lambda}$. In this case, the smallest Lyapunov exponent corresponds to the condition with longest localization length by definition $\gamma=1/\xi$. Therefore, for the localized we expect $\gamma$ is finite; while for the extended phase, we expect that $\gamma L>1$ (or >>1>>1), where $L$ is the chain length. We solve this matrix using the QR decomposition method. In Fig. 3, we plot the Lyapunov exponent $\gamma$ versus $L$ of the quasiperiodic slowly varying model for various $V$ when $E=0$, showing that this value is convergent, which is necessary for a well-defined exponent. These results may also indicate ergodicity, which is necessary during the calculation of the above expression.
@@
To further investigate the localization properties of the quasiperiodic mosaic model with slowly varying potentials in the thermodynamic limit, we plot the Lyapunov exponent $\gamma$ as a function of energy $E$ for various $V$ in Fig. 4. As the energy varies, $\gamma$ changes dramatically from a finite value (indicating localized states) to 0 (indicating extended states), or vice versa, at certain energies. This abrupt jumping in $\gamma$ suggests the presence of MEs in the energy spectrum. When $\kappa=2$, we observe that the Lyapunov exponent at the exact band center is vanishingly small ($\ll 1/L$) for different values of $V$ [see Fig. 4(a)], indicating a resonant state. The MEs are clearly observed around $E_{c}\approx\pm 1.8t$ and $\pm 0.5t$ for $V=0.5t$, and around $E_{c}\approx\pm 1.4t$ for $V=\sqrt{2}t$, which closely align with the exact expression of MEs given by Eq. (6). When $\kappa=3$, we can also find the Lyapunov exponent $\gamma$ for states satisfying $E=\pm t$ is always zero, indicating two resonant states [see Fig. 4(b)]. Additionally, the regime with $\gamma=0$ is consistent with Eqs. (7) and (8).
@@ II.3 Analytical solutions
${\bf u}=(\cdots,u_{i-2},u_{i-1},u_{i},u_{i+1},u_{i+2},\cdots)^{T},$ (19)
with $T$ being the transpose. $H$ is a triangle matrix due to its hopping between nearest neighboring sites. From the equation for $u_{i-1}$ and $u_{i+1}$, where the potentials are zero, we can obtain
@@
Figure 6: Function $\cos(\pi\alpha j^{\nu})$ with (a) $\nu=1.6$ and (b) $\nu=0.6$. (c) Function s=(Ec)Es=(E-c)E with $c>0$ (left panel) and $c<0$ (right panel), and their corresponding regime for $0<s<4t^{2}$, with $t=1$. The red dots denote the MEs.
Next we provide a more detailed explanation of the above results using the method developed by Das Sarma et al. [59, 60] in terms of ”local constancy”. The effective potential $\lambda_{j}$ exhibits the property of becoming locally constant in the thermodynamic limit
@@
due to $0<\nu<1$ [see Fig. 6(b)]. In contrast when $\nu>1$, the oscillating of the potential will become more and more rapid, yielding fully localized states in the thermodynamic limit [see Fig. 6(a)].
@@
with $V\leq c\leq V$, yielding the following simplified equation
@@
The above results have an immediate corollary, that is, the phase boundaries are independent of the phase carried in the potential, assuming $V_{j}=V\cos(\pi\alpha j^{\nu}+\phi)$. The reason is that while this phase will influence the potential at each site, it will not influence the maximal and minimal value of the potential. Meanwhile, the parameters of $\alpha$ is also not important, nor it is not necessary to be an irrational number. We find that $\alpha$ and $\phi$ will never appear in the expression of MEs. These are also verified by our numerical simulation, which are not shown. However, if the phase $\phi$ becomes complex, $\phi=\theta+ih$, the physics may be totally different [64, 22]. To characterize the topological property induced by complex phase, we can define a winding number as
$w(h)=\lim_{L\rightarrow\infty}\frac{1}{2\pi i}\int_{0}^{2\pi}d\theta\frac{\partial}{\partial\theta}\text{log det}[H(\frac{\theta}{L},h)-\epsilon],$ (36)
where we fix the base energy $\epsilon=0$. In Fig. 7(b), we numerically calculate the winding number $w$ as a function of $h$ for the quasiperiodic mosaic slowly varying model with $\kappa=2$ and $V=t$, revealing a topological transition point at $h_{\rm c}\approx 0.42$. To verify whether the topological transition also coincides with a localization-delocalization transition, we present the corresponding fractal dimension $D_{2}$ as a function of $E$ and $V$ when $h=h_{\rm c}$ in Fig. 7(c). It is observed that as $V$ increases to $t$, all states except for the resonant state at $E=0$ become localized, indicating a transition from critical phase to localized phase. In Fig. 7(a), we plot the Lyapunov exponent and its derivative for $V=t$ and $E=t$, which agrees well with the topological transition points $h_{\rm c}\approx 0.42$.
Figure 7: (a) Lyapunov exponent $\gamma$ and its derivative $d\gamma/dE$ and (b) winding number $w$ versus phase $h$ in the quasiperiodic potential $V_{i}=\cos(\pi\alpha i^{\nu}+ih)$. (c) $D_{2}$ of the eigenstates as functions of $E$ and $V$ when $h=h_{\rm c}=0.42t$. Here we can see that the localization-delocalization transition points in (a) and (c) are also topological transition points in (b). Here we choose quasiperiodic mosaic slowly varying models with $\kappa=2$ as an example.